You plan to deposit $2,300 per year for 6 years into a money market account with an annual return of 2%. You plan to make your first deposit one year from today.
Please help me solve this problem. This is the breakdown but I dont understand FVAN=how to get this step
Ordinary annuity:
FVAN = PMT[((1 + I)N ‒ 1)/I]
FVAN = $2,300[((1.02)6 – 1)/0.02]
FVAN = $2,300(6.30812)
FVAN = $14,508.68

Found 2 solutions by greenestamps, ikleyn:
Answer by greenestamps(13216)



From the way you ask the question, I'm not sure what part of this you don't understand.

For one thing, the N in the formula is an exponent:
FVAN = PMT[((1 + I)^N - 1)/I]
FVAN = $2,300[((1.02)^6 – 1)/0.02]

Then the calculations and the answer are correct.

If you are having trouble doing the calculations, it is quite possible that you have some parentheses missing, or in the wrong places. I encourage students to do the calculations one step at a time. For the given numbers....

1+.02 = 1.02
(1.02)^6 = 1.1261624
(1.02)^6-1 = 1.1261624 - 1 = 0.1261624
((1.02)^6-1)/.02 = 0.1261624/.02 = 6.308121
2300*((1.02)^6-1)/.02) = 2300*6.308121 = 14508.68

If your question is about the formula for the future value of an annuity, do an internet search on something like "ordinary annuity formula derivation".

Answer by ikleyn(52936)

.

On Ordinary Annuity,  you have the lessons in this site
    - Ordinary Annuity saving plans and geometric progressions
    - Solved problems on Ordinary Annuity saving plans
in this site.

You can learn  EVERYTHING  related to this subject from these lessons.